OpenStudy (anonymous):

Last year, McBarkie's Burgers contained 200 grams of fat in each burger. This year, the amount of fat has been reduced to 120 grams per burger. What is the percent of decrease in the number of grams of fat per burger to the nearest percent?

OpenStudy (anonymous):

80/200 * 100 = 40% decrease.

OpenStudy (anonymous):

40 %

hero (hero):

The general formula for percent decrease is: \[\%\Delta = \frac{|a - b|}{a} \times 100\] , where a = original amount and b = new amount. In this case, a = 200, b = 120 And 40% decrease is correct.

hero (hero):

Actually, that's the formula for percent increase or decrease

OpenStudy (anonymous):

Yes.